OpenStudy (anonymous):

How does 3 1/2 simplify to 7/2?

OpenStudy (anonymous):

\[\frac{1}{2} = 3+ \frac{1}{2} = \frac{6}{2}+\frac{1}{2}= \frac{7}{2}\] thats the long way, short way: \[a\frac{b}{c} = \frac {ac+b}{c}\] So:\[3\frac{1}{2} = \frac {3(2)+1}{2}= \frac{7}{2}\]

OpenStudy (anonymous):

thank you so much